OpenStudy (anonymous):

Identify the solution(s) of the system of equations, if any. 9x + 2y = 2 8x - 2y = 4

OpenStudy (anonymous):

no need here to multiply any equations here just add both equations together that equals 9x+8x = 2+4 x= 6/17

OpenStudy (anonymous):

ok...I see

OpenStudy (anonymous):

y=-10/17

OpenStudy (anonymous):

(6/17, -10/17)

OpenStudy (anonymous):

So...so let's go back a second and lets talk about how you got to the -10/17 answer.

OpenStudy (anonymous):

ok...I think I see....you did the -4/1(6/17) to equal the -10/17, right?

OpenStudy (anonymous):

u know how to get the x value by elimination method okay ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good then u have the value of x substitute the value of x in any of the two equations u have and u would find the answer of y

OpenStudy (anonymous):

like in this example

OpenStudy (anonymous):

x= 6/17 9x + 2y = 2 soo 9*(6/17)+2y=2

OpenStudy (anonymous):

Oh yeah...duh. I think my mind is slowly shriveling up. LOL

OpenStudy (anonymous):

ha ha I was just typing the same thing to reiterate the answer.

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (anonymous):

it is okay when u have the two values substitute both of them in one of the equations to confirm ur answer ^_^ good luck

OpenStudy (anonymous):

:)